Interest of $840 was earned in 7 months on a balance of $9000. What is the simple interest rate per annum on the account?

Answer by rfer(16322)

I=prt
840=9000*r*7/12
840/9000=7/12r
0.09333=7/12r
0.0933/(7/12)=r
0.0933*12/7=r
0.16=r
r=16%